Rob bought 2 pairs of shorts, 3 shirts and a pair of shoes for $146.64. Jessie bought 3 pairs of shorts, 5 shirts and 2 pairs of shoes for $256.35. Allen bought a pair of shorts and 4 shirts for $104.07. What is the price of a pair of shorts? Express answer rounded to two decimal places.
Answer by mananth(16949)

Rob bought 2 pairs of shorts, 3 shirts and a pair of shoes for $146.64. Jessie bought 3 pairs of shorts, 5 shirts and 2 pairs of shoes for $256.35. Allen bought a pair of shorts and 4 shirts for $104.07. What is the price of a pair of shorts? Express answer rounded to two decimal places.

shorts $x
shirts$y
shoes$z
1.Rob ,.. 2x+3y+z= 146.64
2.Jessie, 3x+5y+2z=256.35
3.Allen , x+4y=104.07
..
multiply eq. 1 by 2 and subtract eq2 from it
4x+6y+2z-3x-5y-2z= 293.28- 256.35
x+y=36.93
..
subtract this equation from eq3
x+4y-x-y=104.07-36.93
3y=67.14
y= 22.38 cost ofshirt.
..
x+y=36.93
x+22.38=36.93
x=36.93-22.38
x=14.55 the cost of shorts.